(i) Calculate divergence A = x²zî - 2y³z²ĵ+ xy²zk at the point (1,-1,1).
1
Expert's answer
2021-10-27T15:36:34-0400
"\dfrac{\partial A_x}{\partial x}=2xz, \dfrac{\partial A_y}{\partial y}=-6y^2z^2,\dfrac{\partial A_z}{\partial z}=xy^2"

"divA=2xz-6y^2z^2+xy^2"

"divA|_{(1,-1,1)}=2(1)(1)-6(-1)^2(1)^2+1(-1)^2"

"=-3"